OpenStudy (anonymous):

Let h(x)=f(g(x)) and p(x)=g(f(x)). Use the table below to compute the following derivatives. a.h'(3) b. p'(2) Anybody here have any remote idea on how to solve this problem?

OpenStudy (anonymous):

Here's the chart

OpenStudy (anonymous):

\[h(x)=f(g(x))\] by the chain rule \[h'(x)=f'(g(x))\times g'(x)\] and so \[h'(3)=f'(g(3))\times g'(3)\] substitute the numbers you know to get what you want

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

yw